# After Retrieval, Before Generation: Enhancing the␣Trustworthiness of Large Language Models in RAG

### v1

@@ 4.1.1 Soft Bias Allocator
Determining soft bias (i.e., $r_{p}$ and $g_{p}$) for a single question presents significant challenges. For instance, the question What is the CPU of iPhone 16? contains no time-sensitive keywords, requiring LLMs to assign bias probabilities through analyzing question semantic and assessing internal knowledge boundaries. To enhance this capability, we leverage TRD’s training data to synthesize annotated examples with analysis paths. We develop two Allocator paradigms: (1) a GRPO-optimized Llama3-8B-Instruct model (pretrained cutoff: Feb 2023), and (2) an ICL-based approach.
@@
Soft Bias Analysis Data Construction. We synthesize data through three steps: (1) We set the retrieval dependency probability $r_{p}^{hard}$ and the generation dependency probability $g_{p}^{hard}$ based on the scenario type in TRD, which can be regarded as the golden label of bias. For TRD’s FA/FI categories, we set $r_{p}^{hard}=0\%$, $g_{p}^{hard}=100\%$; for RA and FE, $g_{p}^{hard}=0\%$, $r_{p}^{hard}=100\%$. (2) Using DeepSeek-R1 and o3-mini as reasoning models, we prompt them to generate analysis path $a$ and $r_{p}^{soft}$ and $g_{p}^{soft}$ (10%-90%) under explicit directional constraints from hard bias labels (e.g., $g_{p}^{soft}>r_{p}^{soft}$ for FA/FI). (3) We select the generated output having a smaller root mean square loss between the soft bias and hard bias labels as high-quality data.
@@
Allocator (GRPO). We employ GRPO to fine-tune our Allocator, using a composite reward function to optimize Llama3-8B-Instruct. We instruct the model to output a tuple $\gamma=(a_{pred},r_{p},g_{p})$, comprising the analysis path $a_{pred}$, predicted retrieval dependency probability $r_{p}$, and generation dependency probability $g_{p}$. The training incorporates four distinct reward components:
Direction Reward: Ensures primary direction alignment between the soft bias and its corresponding hard bias. For question $q$, the direction reward function is defined as:
$R(\gamma,q;direction)=\begin{cases}3&\mathrm{if~{}}\left(r_{p}>g_{p};r_{p}^{hard}>g_{p}^{hard}\right)\mathrm{~{}or~{}}\left(r_{p}<g_{p};r_{p}^{hard}<g_{p}^{hard}\right)\\ 0&\mathrm{otherwise}\end{cases}$ (1)
Format Reward: Verifies output format completeness of $\gamma$ through regular expression matching:
$R(\gamma,q;format)=\left\{\begin{array}[]{ll}1&\mathrm{if~{}}r_{p},g_{p}\mathrm{~{}and~{}}a_{pred}\mathrm{~{}in~{}}\gamma\\ 0&\mathrm{otherwise}\end{array}\right.$ (2)
Sum Reward: Enforces the $r_{p}+g_{p}=100\%$ constraint:
$R(\gamma,q;sum)=\left\{\begin{array}[]{cc}1&\mathrm{if~{}}r_{p}+g_{p}=100\%\\ 0&\mathrm{otherwise}\end{array}\right.$ (3)
Analysis Quality Reward: Promotes semantic similarity between the model’s predicted analysis $a_{pred}$ and high-quality analysis paths generated by larger reasoning models. We compute reward using BGE embeddings to measure textual similarity:
$R(\gamma,q;analysis)=BGE(a_{pred},a)$ (4)
@@
Allocator (ICL). We also implement an in-context learning (ICL) paradigm by retrieving the top-k most similar training examples for each TRD. Instead of relying on soft bias outputs from larger reasoning models, we convert the gold hard-bias presets into soft-bias labels (e.g., 100%→90%, 0%→10%) to construct pseudo soft bias demonstrations. This ensures that BRIDGE’s performance improvements are not attributable to external reasoning model outputs. The analysis is discussed in § 5.2, while implementation details for GRPO and ICL in Allocator are provided in Appendix B.
@@ 4.1.2 Knowledge Collection
Relying solely on the original question for retrieval often yields limited information, prompting recent work to focus on improving retrieval performance in RAG systems to capture broader knowledge weller2022defending . Similarly, generating multiple responses for knowledge-related questions can improve LLM self-consistency and mitigate hallucinations chen2024universal . To expand the knowledge search space, we leverage LLMs to analyze a given question and identify the key information required for resolution. For each key information, the LLM generates a targeted query. These sub-queries act as "probes," enabling both the retriever and the LLM to explore a wider range of knowledge within their respective domains.
Specifically, we allocate sub-queries proportionally to each knowledge source based on the Allocator’s soft bias probabilities. Given $n$ total sub-queries, the number of queries assigned to retrieved knowledge $s_{r}$ and generated knowledge $s_{g}$ is computed as $s_{r}=\lceil r_{p}\times n\rceil$ and $s_{g}=\lceil g_{p}\times n\rceil$, respectively. The first $s_{r}$ sub-queries are processed by the Multi-query Retriever, aggregating external documents into the retrieved knowledge $K_{ret}$, while the first $s_{g}$ sub-queries are answered by the Multi-query Generator, forming the generated knowledge $K_{gen}$. This adaptive allocation improves the effectiveness of knowledge searching.
@@ 4.2.1 Knowledge Scorer
According chen2024bge , our BGE-based Scorer computes three similarity metrics between knowledge sources: sparse matching for lexical overlap, dense matching for semantic similarity, and ColBERT khattab2020colbert matching for token-level relevance, respectively weighted as 0.2, 0.4, and 0.4 by grid search in the TRD validation set. This multi-granular approach better assesses knowledge relevance.
@@ 4.2.2 Maximum Soft-bias Decision Tree
To determine the optimal integration strategy, we develop a Maximum Soft-bias Decision Tree that evaluates knowledge consistency through four knowledge matching scores and soft bias to help LLM make the final decision. As shown in Figure 2, we get four matching scores from Scorer. Based on these scores and soft bias, the tree executes three sequential checks:
@@
Conflict Detection. Trigger FA if $S_{1}+S_{4}>(1-S_{1})+(1-S_{4})$ (implying high consistency in two knowledge sources), the LLM will integrate all knowledge.
Confidence Calculation. We compute the trustworthiness score for the retriever and the LLM:
$T_{Ret}=r_{p}[S_{3}+(1-S_{2})]$ (5)
$T_{LLM}=g_{p}[S_{2}+(1-S_{3})]$ (6)
@@
This formulation reflects two key insights: (1) Higher dependency probability ($r_{p}$ or $g_{p}$) amplifies the impact of corresponding knowledge matches, and (2) A mismatch between knowledge from one source reduces the trustworthiness score, while a match between knowledge from another knowledge source also decreases its score. For example, poor alignment between generated knowledge and internal knowledge ($S_{2}\rightarrow 0$), along with good alignment between retrieved knowledge and external knowledge ($S_{3}\rightarrow 1$), both contribute to lowering $T_{LLM}$.
@@
Threshold-based Decision. As shown in Figure 2, we employ two thresholds ($\alpha=0.5$, $\beta=1.1$) determined through grid search on TRD’s validation set. For cases where $T_{Ret}<T_{LLM}$, the system adopts FI when $T_{LLM}>\alpha$, otherwise triggering RA. When $T_{Ret}\geq T_{LLM}$, three outcomes emerge: (1) FE if $T_{Ret}\geq\beta$, (2) RA if $T_{Ret}<\alpha$, or (3) Reflection mechanism adjusts the sub-queries for intermediate cases ($\alpha\leq T_{Ret}<\beta$). After three unsuccessful cycles, the system defaults to RA as a conservative response. The asymmetric thresholds ($\beta>\alpha$) mean the retriever knowledge source requires higher trustworthiness ($\beta=1.1$) due to potential retrieval noise.
Finally, our framework dynamically adjusts response strategies based on predicted decisions: FA employs all knowledge; FI relies solely on $K_{int}$ and $K_{gen}$; FE prioritizes $K_{ext}$ and $K_{ret}$; and RA responds with "I don’t know". The theoretical analysis of GRPO, the hyper-parameter settings, and all system prompts in BRIDGE are detailed in Appendix B.
@@ 5.1 Experimental Setups
Datasets. We evaluated all methods on four datasets. (1) TRD: Our constructed test set. (2) RealtimeQA (RQA) kasai2024realtime : A dynamic news QA platform. We select 468 questions emerging after October 2023, providing their corresponding evidence as external knowledge. This dataset aligns with the FE scenario. (3) HotpotQAPoisoned (HQAP) yang2018hotpotqa : We use the version from RobustRAG xiang2024certifiably , where external knowledge is corrupted via PoisonedRAG zou2024poisonedrag . Thus, we generate valid internal knowledge from correct answers using the method in § 3.2. This dataset corresponds to the FI scenario. (4) ConflictBank su2024conflictbank : A multiple-choice dataset containing temporal misalignment facts fabricated from Wikidata entries with future timestamps. We use the original correct evidence as internal knowledge and fabricated facts as external knowledge, and randomly sample 1,000 questions to evaluate LLMs under the FI scenario.
@@
RAG settings & Baselines. (1) RAG settings: We evaluate RAG baselines using three LLMs (GPT-3.5-turbo, Qwen 72B, and Llama3-8B-Instruct) with parametric knowledge cutoffs before February 2023. To ensure fairness, all methods employ identical retriever parameters and are granted access to the English Wikipedia dump (cutoff date: February 1, 2025). (2) Baselines: Our baselines contain various aspects of LLM trustworthiness in RAG systems. Prompt-based methods include OPIN zhou2023context , USC chen2024universal , RobustRAG xiang2024certifiably , InstructRAGICL wei2025instructrag , AstuteRAG wang2024astute , and TrustRAG zhou2025trustrag . GPT does not support deterministic decoding; hence, we report results averaged from three runs. Fine-tuning-based methods encompass SelfRAG asai2023self , CAD shi2024trusting , InstructRAGFT wei2025instructrag and Trust-Align song2025measuring . Detailed settings are available in Appendix C.
5.2 Experiment Results
@@
Main Results. As shown in Table 1, compared to vanilla RAG, OPIN, which prioritizes external knowledge, achieves notable gains on RQA, but suffers performance drops on HQAP and ConflictBank. Conversely, USC, which enforces internal knowledge consistency, excels on HQAP and ConflictBank but deteriorates sharply on RQA. Knowledge integration methods (RobustRAG, InstructRAG, AstuteRAG, and TrustRAG) demonstrate consistent improvements over vanilla RAG on TRD, underscoring the effectiveness of combining different knowledge sources. However, their lower RR scores indicate that their response strategies for refusing to answer are insufficiently considered.
@@
CAD, which lacks mechanisms to resolve knowledge conflicts or reject uncertain questions, underperforms across most scenarios. In contrast, SelfRAG, InstructRAGFT and Trust-Align, which dynamically select knowledge sources or validate knowledge boundaries, achieve significantly higher RR rates. Nevertheless, their simple knowledge integration strategy limits broader applicability in all scenarios. As described in §4.1.1, BRIDGEICL, despite relying solely on templated soft bias demonstrations (without using data from large-scale reasoning models), outperforms baselines, validating benefits stem from soft bias design rather than synthetic data from larger reasoning models. BRIDGEGRPO further excels with its adaptive bias allocation, achieving balanced performance across all scenarios. Notably, BRIDGE has a sum of 100% for Acc and RR on the ConflictBank dataset when using GPT-3.5-Turbo and Qwen72B. This demonstrates our model’s highly reliable response strategy: when unable to provide correct answers using its internal knowledge, it consistently chooses to abstain from responding rather than risking incorrect outputs.
Table 1: Overall results, reporting accuracy (Acc), refusal rate (RR, regardless of question answerability), and exact match score (EM) rajpurkar2016squad , with the best results highlighted in bold.
Methods TRD RQA HQAP ConflictBank
Acc(%) RR(%) EM(%) EM(%) Acc(%) RR(%)
Based on GPT-3.5-turbo
Vanilla 53.71 5.81 72.50 44.40 52.04 0.20
OPIN zhou2023context 63.93 3.34 76.25 23.47 48.03 2.00
USC chen2024universal 65.53 3.01 26.25 70.70 73.46 1.08
RobustRAG xiang2024certifiably 62.50 6.69 65.82 64.00 58.40 3.04
InstructRAGICL wei2025instructrag 65.55 4.29 74.36 57.30 57.45 1.71
AstuteRAG wang2024astute 64.26 2.81 74.03 43.43 57.45 1.66
TrustRAG zhou2025trustrag 66.39 3.42 78.48 40.00 52.01 2.50
BRIDGEICL 72.32 31.40 76.92 70.97 61.86 38.14
BRIDGEGRPO 77.90 33.33 78.48 72.16 75.31 24.69
Based on Qwen 72B
Vanilla 63.19 9.50 70.25 61.61 57.83 0.61
OPIN zhou2023context 63.40 3.41 75.95 32.65 42.22 4.44
USC chen2024universal 64.20 4.80 28.75 76.00 70.74 2.04
RobustRAG xiang2024certifiably 64.33 7.28 74.25 65.00 59.72 3.04
InstructRAGICL wei2025instructrag 64.69 4.71 73.75 67.32 56.80 1.08
AstuteRAG wang2024astute 66.29 3.60 75.00 66.35 66.89 1.73
TrustRAG zhou2025trustrag 66.41 3.25 74.68 67.47 63.75 1.25
BRIDGEICL 74.20 39.56 76.25 72.15 67.35 32.65
BRIDGEGRPO 85.48 32.34 84.06 75.00 69.39 30.61
Based on Llama3 8B-Instruct
Vanilla 42.20 1.80 68.75 57.57 46.00 0.81
SelfRAG asai2023self 47.01 14.97 61.25 55.42 61.25 8.18
CAD shi2024trusting 24.40 1.50 55.00 35.00 39.20 1.95
InstructRAGFT wei2025instructrag 51.81 13.43 69.30 70.70 60.78 15.00
Trust-Align song2025measuring 53.05 18.11 72.50 69.30 65.61 5.10
BRIDGEICL 62.94 23.86 78.75 71.72 59.60 31.46
BRIDGEGRPO 73.33 19.30 80.00 78.72 71.72 28.28
Detailed Results. As shown in Figure 3, we compare the response accuracy of different methods across various RAG scenarios in TRD. BRIDGE demonstrates balanced capabilities across all scenarios, achieving performance comparable to state-of-the-art baselines on FA, FI, and FE cases while attaining optimal performance on RA scenarios. Figure 4 presents the error response strategy distribution of BRIDGE based on GPT-3.5-Turbo. The results reveal that FA exhibits the lowest error probability, whereas FI shows the highest. Notably, most FI cases are misclassified as FA, preserving the model’s ability to access the internal knowledge source. For FE misclassifications, the model typically adopts a conservative rejection strategy to maintain credibility. The misclassification of RA as FE represents an understandable scenario, as current RAG baselines tend to favor incorrect external knowledge over refusing to answer. Our method has minimized such potential harm through its carefully designed response mechanism.

### v2

@@ 3.1.1 Soft Bias Allocator
Determining soft bias (i.e., $r_{p}$ and $g_{p}$) for a single question presents significant challenges. For instance, the question What is the CPU of the iPhone 16? contains no time-sensitive keywords, requiring the LLM to assign dependency probabilities by analyzing the question’s semantics and assessing its internal knowledge boundaries. To enhance this capability, we leverage TRD’s training data to construct annotated examples with analysis paths. We develop two Allocator paradigms: (1) a GRPO-optimized Llama3-8B-Instruct (pretrained cutoff: Feb 2023), and (2) an ICL-based approach.
@@
Soft Bias Analysis Data Construction. We construct data through three steps: (1) We assign the retrieval dependency probability $r_{p}^{hard}$ and the generation dependency probability $g_{p}^{hard}$ according to the scenario type in TRD. These hard bias values serve as the ground-truth labels for soft bias learning. For TRD’s FA/FI categories, we set $r_{p}^{hard}=0\%$, $g_{p}^{hard}=100\%$; for RA and FE, $g_{p}^{hard}=0\%$, $r_{p}^{hard}=100\%$. (2) Using DeepSeek-R1 and o3-mini as reasoning models, we instruct them to generate an analysis path $a$, along with soft bias probabilities $r_{p}^{soft}$ and $g_{p}^{soft}$ constrained to the range $10\%,90\%]$, under explicit directional guidance derived from the hard bias labels (e.g., $g_{p}^{soft}>r_{p}^{soft}$ for FA/FI queries). (3) We filter for high-quality data by selecting the generated outputs that exhibit the lowest root mean squared error between the soft ($r_{p}^{soft},g_{p}^{soft}$) and hard ($r_{p}^{hard},g_{p}^{hard}$) bias labels.
@@
Allocator (GRPO). We employ GRPO to fine-tune our Allocator, using a composite reward function to optimize Llama3-8B-Instruct. We instruct the model to output a tuple $\gamma=(a_{pred},r_{p}^{pred},g_{p}^{pred})$, comprising the analysis path $a_{pred}$, predicted retrieval dependency probability $r_{p}^{pred}$, and generation dependency probability $g_{p}^{pred}$. The training objective consists of four distinct reward components: Direction Reward, Format Reward, Sum Reward, and Analysis Quality Reward. The specific functions and designs of these rewards are detailed in the Appendix B.1.2.
@@
Allocator (ICL). We also implement an ICL paradigm by retrieving the top-k similar training examples for each TRD. Instead of relying on soft-bias outputs and analysis paths from reasoning models, we directly convert the gold hard-bias presets into soft-bias labels (e.g., 100%→90%, 0%→10%) to construct pseudo soft-bias demonstrations. This ensures that BRIDGE’s performance improvements are not attributable to external reasoning model abilities. We discuss the analysis in § 4.2 and provide implementation details of the Allocator (GRPO and ICL) in Appendix B.1.
@@ 3.1.2 Knowledge Collection
[…]
@@ 3.2.1 Knowledge Scorer
In this stage, we compute three similarity measures between knowledge sources using the BGE Scorer Chen et al. (2024a): sparse matching to capture lexical overlap, dense matching to assess semantic similarity, and ColBERT-based matching Khattab and Zaharia (2020) to capture token-level correlations. The respective weights are determined by grid search on the TRD validation set (details are provided in Appendix B.2). This multi-granularity strategy provides a more comprehensive evaluation of knowledge relevance and demonstrates broad applicability. As validated in § 4.2, all weights and hyperparameters, tuned exclusively on TRD, achieve balanced performance when applied directly to other datasets without further modification, thereby confirming their robustness.
@@ 3.2.2 Maximum Soft-bias Decision Tree
After obtaining four matching scores using the Scorer, we perform three sequential checks:
@@
Consistency Detection. If $S_{1}+S_{4}>(1-S_{1})+(1-S_{4})$ (implying high consistency between the two knowledge sources), the system triggers FA, and the LLM integrates all knowledge.
@@
Confidence Calculation. We compute the trustworthiness scores for the retriever and the LLM as follows: $T_{Ret}=r_{p}^{pred}[S_{3}+(1-S_{2})]$, $T_{LLM}=g_{p}^{pred}[S_{2}+(1-S_{3})]$. This formulation reflects two key insights: (1) a higher dependency probability ($r_{p}^{pred}$ or $g_{p}^{pred}$) amplifies the influence of the corresponding knowledge matches, and (2) inconsistencies in one knowledge source reduce its trustworthiness score, while consistency in the other source may further attenuate the first source’s score. For example, poor alignment between $K_{gen}$ and $K_{int}$ ($S_{2}$→0) and good alignment between $K_{ret}$ and $K_{ext}$ ($S_{3}$→1) both contribute to lowering $T_{LLM}$.
@@
Threshold-based Decision. As shown in Figure 2, we employ two thresholds ($\alpha$, $\beta$) determined by grid search on the TRD validation set (see Appendix B.2). For cases where $T_{Ret}<T_{LLM}$, the system adopts FI when $T_{LLM}>\alpha$, otherwise triggering RA. When $T_{Ret}\geq T_{LLM}$, the system’s response is determined by the value of $T_{Ret}$: it triggers FE if $T_{Ret}\geq\beta$, initiates RA if $T_{Ret}<\alpha$, or invokes the Reflection mechanism Madaan et al. (2023) to adjust sub-queries for intermediate cases ($\alpha\leq T_{Ret}<\beta$). After three unsuccessful cycles, the system defaults to RA as a conservative response. The asymmetric thresholds ($\beta>\alpha$) mean the retriever knowledge source requires higher trustworthiness due to potential retrieval noise. Finally, our framework dynamically adjusts response strategies based on predicted interpretable decisions: FA employs all knowledge; FI relies solely on $K_{int}$ and $K_{gen}$; FE prioritizes $K_{ext}$ and $K_{ret}$; and RA responds with "I don’t know". All system prompts are detailed in Appendix B.3.
Methods TRD RQA HQAP ConflictBank
Simu Real Simu Real Simu Real Simu Real
Acc(%) RR(%) Acc(%) RR(%) EM(%) EM(%) EM(%) EM(%) Acc(%) RR(%) Acc(%) RR(%)
Based on GPT-3.5-turbo
Vanilla 53.71 5.81 55.80 14.61 72.52 29.41 44.40 69.23 52.04 0.20 38.81 38.24
OPIN Zhou et al. (2023) 63.93 3.34 54.62 16.44 76.25 23.75 23.47 54.55 48.03 2.00 37.02 40.71
USC Chen et al. (2024c) 65.53 3.01 50.41 15.77 26.25 25.71 70.70 73.47 73.46 1.08 40.21 29.87
RobustRAG Xiang et al. (2024) 62.50 6.69 56.94 14.69 65.82 29.11 64.00 73.00 58.40 3.04 39.39 25.25
InstructRAGICL Wei et al. (2025) 65.55 4.29 57.11 20.84 74.36 28.75 57.30 54.55 57.45 1.71 37.71 40.02
AstuteRAG Wang et al. (2024) 64.26 2.81 58.60 14.42 74.03 26.25 43.43 71.12 57.45 1.66 31.15 46.28
TrustRAG Zhou et al. (2025) 66.39 3.42 55.53 15.03 78.48 32.50 40.00 74.85 52.01 2.50 27.66 42.55
BRIDGEICL 72.32 31.40 68.72 32.43 76.92 33.66 70.97 79.69 61.86 38.14 76.15 28.42
BRIDGEGRPO 77.90 33.33 75.66 22.71 78.48 33.38 72.16 81.23 75.31 24.69 77.14 24.61
Based on Qwen 72B
Vanilla 63.19 9.50 52.20 14.82 70.25 30.12 61.61 71.31 57.83 0.61 51.37 31.72
OPIN Zhou et al. (2023) 63.40 3.41 58.22 54.53 75.95 31.22 32.65 59.21 42.22 4.44 23.02 57.21
USC Chen et al. (2024c) 64.20 4.80 53.51 8.82 28.75 26.25 76.00 86.97 70.74 2.04 62.41 16.32
RobustRAG Xiang et al. (2024) 64.33 7.28 54.07 9.76 74.25 26.25 65.00 72.81 59.72 3.04 55.41 24.71
InstructRAGICL Wei et al. (2025) 64.69 4.71 53.27 9.66 73.75 25.17 67.32 72.61 56.80 1.08 53.16 17.48
AstuteRAG Wang et al. (2024) 66.29 3.60 55.83 7.61 75.00 23.08 66.35 86.62 66.89 1.73 52.17 19.26
TrustRAG Zhou et al. (2025) 66.41 3.25 54.31 5.41 74.68 24.05 67.47 83.88 63.75 1.25 53.81 16.48
BRIDGEICL 74.20 39.56 58.78 27.32 76.25 32.11 72.15 85.32 67.35 32.65 74.18 23.81
BRIDGEGRPO 85.48 32.34 76.86 27.07 84.06 36.71 75.00 87.61 69.39 30.61 75.27 16.82
Based on Llama3 8B-Instruct
Vanilla 42.20 1.80 51.80 19.80 68.75 25.00 57.57 64.00 46.00 0.81 33.00 22.00
SelfRAG Asai et al. (2023) 47.01 14.97 48.92 18.26 61.25 27.12 55.42 68.31 61.25 8.18 36.25 15.21
CAD Shi et al. (2024) 24.40 1.50 45.21 19.15 55.00 20.00 35.00 58.13 39.20 1.95 20.13 27.75
InstructRAGFT Wei et al. (2025) 51.81 13.43 53.61 15.32 69.30 23.75 70.70 41.07 60.78 15.00 54.28 16.07
Trust-Align Song et al. (2025) 53.05 18.11 48.62 21.61 72.50 16.25 69.30 33.00 65.61 5.10 20.80 21.45
BRIDGEICL 62.94 23.86 61.52 33.67 78.75 32.52 71.72 73.11 59.60 31.46 41.21 21.66
BRIDGEGRPO 73.33 19.30 72.02 23.68 80.00 32.91 78.72 90.42 71.72 28.28 62.42 18.64
Table 1: Results are reported in terms of accuracy (Acc), rejection rate ($R$, regardless of whether the question is answerable), and exact match score (EM) Rajpurkar et al. (2016), with the best results highlighted in bold.
@@ 4.1 Experimental Setups
Datasets. We evaluate all methods on four datasets. (1) TRD: This is the test set that we constructed. (2) RealtimeQA (RQA) Kasai et al. (2024): A dynamic news QA platform. We select 468 questions that emerged after October 2023 and provide their corresponding evidence as external knowledge. This dataset aligns with the FE scenario. (3) HotpotQAPoisoned (HQAP)Yang et al. (2018): We use the version from RobustRAG Xiang et al. (2024), where external knowledge is corrupted by PoisonedRAG Zou et al. (2024). This dataset corresponds to the FI scenario. (4) ConflictBankSu et al. (2024): This is a multiple-choice dataset containing temporally misaligned external knowledge fabricated from Wikidata entries with future timestamps. We designate the fabricated temporal misalignment entries as external knowledge. Then, we randomly sample 1,000 questions from the dataset to evaluate the baselines under the FI scenario. The internal knowledge generation for all LLMs follows § 2.2.
@@
Baselines. Our baselines contain various aspects of LLM trustworthiness in RAG systems. Prompt-based methods include OPIN Zhou et al. (2023), USC Chen et al. (2024c), RobustRAG Xiang et al. (2024), InstructRAGICL Wei et al. (2025), AstuteRAG Wang et al. (2024), and TrustRAG Zhou et al. (2025). GPT does not support deterministic decoding; hence, we report results averaged from three runs. Fine-tuning-based methods encompass SelfRAG Asai et al. (2023), CAD Shi et al. (2024), InstructRAGFT Wei et al. (2025) and Trust-Align Song et al. (2025).
RAG Setup. We evaluate RAG baselines using three backbone LLMs with knowledge cutoff dates prior to February 2023. For rigorous quantitative assessment, we first conduct experiments in a controlled setting where external knowledge is explicitly provided, simulating the initial output of a retriever; this setup is referred to as Simu. To further assess the practical applicability of these methods, we implement a realistic scenario termed Real, in which the retriever queries a Wikipedia dump using the user’s question and dynamically retrieves the top-3 documents as external knowledge $K_{\text{ext}}$. To ensure a fair comparison, all methods employ the same retriever configuration and access identical Wikipedia dumps, with a knowledge cutoff date of February 1, 2025. Further details on the RAG setup are provided in Appendix C.
4.2 Main Results
@@
As shown in Table 1, compared to vanilla RAG, OPIN, which prioritizes external knowledge, achieves notable gains on RQA, but suffers performance drops on HQAP and ConflictBank. Conversely, USC, which enforces internal knowledge consistency, excels on HQAP and ConflictBank but deteriorates sharply on RQA. Knowledge integration methods (RobustRAG, InstructRAG, AstuteRAG, and TrustRAG) demonstrate consistent improvements over vanilla RAG on TRD, underscoring the effectiveness of combining different knowledge sources. However, their lower RR scores indicate that their response strategies for refusing to answer are insufficiently considered.
@@
CAD, which lacks mechanisms to resolve knowledge conflicts or reject uncertain questions, underperforms across most scenarios. In contrast, SelfRAG, InstructRAGFT, and Trust-Align, which dynamically select knowledge sources or validate knowledge boundaries, achieve significantly higher RR rates. Nevertheless, their simple knowledge integration strategy limits broader applicability in all scenarios. As described in §3.1.1, BRIDGEICL, despite relying solely on templated soft bias demonstrations (without using data from large-scale reasoning models), outperforms baselines, which validates that the benefits arise from the soft-bias design rather than from synthetic data produced by larger reasoning models. BRIDGEGRPO further excels with its adaptive bias allocation, achieving balanced performance across all scenarios. Notably, on the ConflictBank dataset, BRIDGE achieves a combined Acc and RR of 100% when using GPT-3.5-Turbo and Qwen72B. This demonstrates our model’s highly reliable response strategy: when unable to provide correct answers using its internal knowledge, it chooses to abstain answering rather than risking incorrect outputs.
4.3 Detailed Results
As shown in Figure 3, we compare the response accuracy of different methods across various RAG scenarios in TRD. BRIDGE demonstrates balanced capabilities across all scenarios, achieving performance comparable to state-of-the-art baselines on FA, FI, and FE cases while attaining optimal performance on RA scenarios.